OpenStudy (anonymous):

what is the sum of the geometric sequence -4, 12, -36 if there are 7 terms

OpenStudy (ghazi):

s= n/2(2*a+(n-1)d) apply this

OpenStudy (anonymous):

i dont understand how to apply can u show me in steps? :

jimthompson5910 (jim_thompson5910):

wrong formula ghazi, that formula applies to arithmetic sequences

OpenStudy (anonymous):

i need geometric sequences

OpenStudy (cwrw238):

sum of n terms = a * r^n - 1 ------ r - 1 where a = first term and r = common ratio common ratio for this series is 12 / -4 = -3 and a = -4

OpenStudy (ghazi):

oh yea....sorry that was a mistake

OpenStudy (anonymous):

if the 7th term is in a geometric sequence is 64 and the common ratio is -2 what are the 1st and 14th term ?

OpenStudy (amistre64):

please post new questions to the left after closing old ones.

OpenStudy (cwrw238):

n = 7 sum 7 terms = -4 * (-3)^2 - 1 --------- -3-1

OpenStudy (cwrw238):

sorry - replace (-3)^2 with (-3)^7 (n=7)

OpenStudy (anonymous):

so the answere is n=7 ?

OpenStudy (cwrw238):

no - 7 is number of terms sum = (-4) * (-3)^7 - 1 --------- (-4) - calculator job

OpenStudy (anonymous):

so what would be the answere?

OpenStudy (cwrw238):

note (-4) cancel out so its (-3)^7 - 1 - you can do this on a calculator

OpenStudy (cwrw238):

- can't you?

OpenStudy (anonymous):

-21888

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

???

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

thankuu so much! tehre is one more question i have if that wouldnt bother u to help please??